M(oH)3 equals 78 g/mol what the atomic mass


1
Expert's answer
2021-11-15T10:54:54-0500

Solution:

Mr(M(OH)3) = 78 g/mol

Mr(M(OH)3) = Ar(M) + 3×Ar(O) + 3×Ar(H)


Ar(H) = 1 g/mol

Ar(O) = 16 g/mol


Therefore,

Ar(M) = Mr(M(OH)3) - 3×Ar(O) - 3×Ar(H)

Ar(M) = 78 - 3×16 - 3×1 = 27 (g/mol)

Ar(M) = 27 g/mol - aluminium (Al)


Answer: The atomic mass of M is 27 g/mol